OpenStudy (aroub):

John's boat travels at a speed of 20km/hr in still water. He covers a distance of 90 km rowing against the current. He then covers the same distance rowing with the current. The total time in the two cases is 9 hours and 36 minutes. Find the speed of the current. Is it 5 ?

OpenStudy (ash2326):

Let speed of river be \[ \frac{90}{20-5}+\frac{90}{20+5}=6+3.6= 9\ hours\ 36\ minutes\] so you're correct:D

OpenStudy (aroub):

Oh great! I finally got something correct :) They're very hard :( But there are many ways to solve this problem right? -Thank you so much!! :D

OpenStudy (ash2326):

You're welcome, these are easy . We just gotta keep some points in mind:)

OpenStudy (ash2326):

If you have any doubt. I'd be glad to help

OpenStudy (aroub):

Yes I know! But.. I have some difficulties solving these problems. Hehe, I hope I don't find any, but if I found then THANK YOU :D

OpenStudy (ash2326):

I also hope so. Best of luck:)

OpenStudy (aroub):

TY!